Let f(x)=3x^2-2 and g(x)=4x+4 Find each value.
(g o f)(√3)

Answer by CubeyThePenguin(3113)

(g o f)(sqrt(3))
= g(f(sqrt(3))
= g(3(sqrt(3))^2 - 2)
= g(7)
= 4(7) + 4
= 32